OpenStudy (anonymous):

will fan and medal The surface area of a triangular pyramid is 375.2 m2. What would change if the surface area were measured in square centimeters? A. The surface area of the pyramid would decrease. B. The surface area of the pyramid would increase. C. The surface area of the pyramid would stay the same, but the number representing the surface area would decrease. D. The surface area of the pyramid would stay the same, but the number representing the surface area would increase.

OpenStudy (kc_kennylau):

Do you know how to convert from m^2 to cm^2?

OpenStudy (anonymous):

1 m= 100 cm

OpenStudy (kc_kennylau):

That's only m to cm, what about m^2 to cm^2?

OpenStudy (anonymous):

idk

OpenStudy (kc_kennylau):



OpenStudy (kc_kennylau):

What's the area of this square in m^2?

OpenStudy (kc_kennylau):

And the area of this square in cm^2?

OpenStudy (kc_kennylau):

Do you know how to calculate the area of a square?

OpenStudy (anonymous):

lxw

OpenStudy (anonymous):

1m=100cm 1m^2=100^2 1m^2=10000 if you get first area 6m^2 2th one is 6 * 10000cm^2 the number is more but area is first one ;)

OpenStudy (kc_kennylau):

Then how did you get 40000cm^2?

OpenStudy (kc_kennylau):

Never mind, I just wanted to prove that, but I think I should skip the proof

OpenStudy (kc_kennylau):

So we know that 1m^2 = 10000cm^2

OpenStudy (kc_kennylau):

Please express the surface area of the pyramid in terms of cm^2